OpenStudy (anonymous):

(x - c)/2 = d solve for x. I got x = 2d + c. Is that correct?

OpenStudy (anonymous):

First off, I would put both in fraction form since one is already. \[\frac{ (x-c) }{ 2 } = \frac{ d }{ 1 }\] Now you want to make both fractions equal so they can cancel out. \[\frac{ (x-c) }{ 2 } = \frac{ 2d }{ 2 }\] Now the new formula is: x-c = 2d Isolate x from here.

OpenStudy (anonymous):

I see what you are doing. I follwed these steps and ended up with : x = 2d + c. So this is the correct answer then?

OpenStudy (anonymous):

Yes, that's correct.

OpenStudy (anonymous):

woo hoo! thank you!!

OpenStudy (anonymous):

No Problem, Glad I could help ;)